OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (anonymous):

cos(x) = sin(x), and you want to find all the solutions within the given interval of [0, 2π). Just remember that we can re-write cos(x) = sin(x) as: cos(x) = sin(x) tan(x) = 1 Reason why: sin(x) / cos(x) = tan(x) And as a result: tan(x) = 1 where x = π/4 and 7π/4

OpenStudy (harsimran_hs4):